OpenStudy (anonymous):

1/6 (3/4x-2) = -1/5 solving for x

OpenStudy (anonymous):

(1/8)x-1/3=-1/5 1/8x=-1/5+1/3 @lolmath ,can u continue?

OpenStudy (anonymous):

@lolmath ,??

OpenStudy (anonymous):

i dont understand how you got 1/8

OpenStudy (anonymous):

1/6*3/4=1/8

OpenStudy (anonymous):

(1/6*3/4)*x-1/6*-2=1/8*x-1/3

OpenStudy (anonymous):

@lolmath ,did u get it?

OpenStudy (anonymous):

oh ok i got it. and i cant do anything after i move 1/3 over right? so the answer is what you stated as the last equation

OpenStudy (anonymous):

1/8x=-1/5+1/3

OpenStudy (anonymous):

yes,right

OpenStudy (anonymous):

-1/5+1/3=2/15 is it ok?

OpenStudy (anonymous):

it seems to me like the answer would be 1/8x=2/15

OpenStudy (anonymous):

now u have to do this: 2/15*8=16/15

OpenStudy (anonymous):

so x=16/15

OpenStudy (anonymous):

@lolmath ,do u understand?

OpenStudy (anonymous):

i do! thank you. the fractions were a bit overwhelming.

OpenStudy (anonymous):

:)